OpenStudy (anonymous):

Hi, I've got the Barron's ACT Flashcard Set and one of the questions seems a little off: Multiple-Choice 76 If (3x + a) · (x + b) = 3x^2 - 19x - 14, what are a and b? A. a=1, b=19 B. 1=2, b=7 C. a=-1, b=14 D. a=2, b=-7 E. a=1, b=-14 I did the FOIL, but it doesn't get rid of x all the way, so I can't figure out how to get a and b. Thank you!

OpenStudy (welshfella):

first expand (3x + a)(x + b) = 3x^2 + 3bx + ax + ab = 3x^2 + x(3b + a) + ab

OpenStudy (welshfella):

now compare coefficients with 3x^2 - 19x - 14 = 0

OpenStudy (welshfella):

this gives 3b + a = -19 ab = -14 solve this system of equations

OpenStudy (welshfella):

* ignore the '= 0' call it a typo!

OpenStudy (anonymous):

What does it mean to compare coefficients? Wouldn't that be to set them equal to each other?

OpenStudy (welshfella):

yes the coefficient of x in one expression is 3a+b and the other its -19 so we can set them equal

OpenStudy (welshfella):

i should have said equate coefficents

OpenStudy (anonymous):

So wouldn't that give us 3x^2 + x(3b + a) + ab = 3x^2 - 19x - 14 ?

OpenStudy (welshfella):

right - though there should be an identity sign not an equal sign in between

OpenStudy (welshfella):



OpenStudy (welshfella):

an identity differs from an equation in that an identity is true for all values of the variable as its an identity we can equate coefficients

OpenStudy (welshfella):

Strictly speaking the original equation should have used the identity sign not the '='.